OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

so the method we will use is just substitution so you say since x - y = 16, x = 16 + y then plug that into 2x + y = 2 so you get 2(16 + y) + y = 2 32 + 2y + y = 2 32 + 3y = 2 so, 3y = -30 y = -10 then we can go back an plug in that y value to solve for x so x - (-10) = 16 so x = 6